# Decentralized Optimization with Coupled Constraints

### v3

@@ 5.1 Algorithm
As stated in Lemma 4, problem (21) is equivalent to problem (1). Due to Lemma 1, its objective is strongly convex, allowing us to apply Algorithm 1 to it. Using Lemma 3, we obtain that the condition numbers of $\mathbf{W^{\prime}}$ and $\mathbf{K}$ are bounded as $O(1)$, but a single multiplication by $\mathbf{W^{\prime}}$ and $\mathbf{K},\mathbf{K}^{\top}$ translates to $O(\sqrt{\kappa_{\mathbf{W}}})$ multiplications by $\mathbf{W}$ and $O(\sqrt{\kappa_{\mathbf{B}}})$ multiplications by $\mathbf{B},\mathbf{B}^{\top}$ respectively.
We implement multiplications by $\mathbf{W^{\prime}}$ and $\mathbf{K},\mathbf{K}^{\top}$ through numerically stable Chebyshev iteration procedures given in Algorithms 3 and 5, which only use decentralized communications and multiplications by $\mathbf{A},\mathbf{A}^{\top}$. Lemmas 1, 2 and 3 allow us to express the complexity of Algorithm 1 in terms of the parameters of the initial problem given in Assumptions 1, 2 and 3. All this leads us to the following Theorem 1, a detailed proof of which is provided in Appendix A.3, as well as the derivation of Algorithms 3 and 5 and values of the parameters of Algorithm 2.
@@
2:$\rho\coloneqq\left(\sqrt{L_{\mathbf{W}}}-\sqrt{\mu_{\mathbf{W}}}\right)^{2}/16$
3:$\nu\coloneqq\left(\sqrt{L_{\mathbf{W}}}+\sqrt{\mu_{\mathbf{W}}}\right)/2$
4:$\delta^{0}\coloneqq-\nu/2$,  $n\coloneqq\lceil\sqrt{\kappa_{\mathbf{W}}}\rceil$
5:$p^{0}\coloneqq-\mathbf{W}y/\nu$,  $y^{1}\coloneqq y+p^{0}$
@@
2:$z\coloneqq r\left(\mathbf{A}x+\gamma\cdot\mathrm{\mathbf{mul}\mathbf{W^{\prime}}}(y)-\mathbf{b}\right)$
@@
2:$\rho\coloneqq\left({L_{\mathbf{B}}}-{\mu_{\mathbf{B}}}\right)^{2}/16$
3:$\nu\coloneqq\left({L_{\mathbf{B}}}+{\mu_{\mathbf{B}}}\right)/2$
4:$\delta^{0}\coloneqq-\nu/2$,  $n\coloneqq\lceil\sqrt{\kappa_{\mathbf{B}}}\rceil$
5:$q^{0}\coloneqq\mathbf{A}x+\gamma\cdot\mathrm{\mathbf{mul}\mathbf{W^{\prime}}}(y)-\mathbf{b}$
6:$p^{0}\coloneqq-\dfrac{1}{\nu}\begin{pmatrix}\mathbf{A}^{\top}q^{0}\\ \gamma\cdot\mathrm{\mathbf{mul}\mathbf{W^{\prime}}}(q^{0})\end{pmatrix}$
7:$u^{1}\coloneqq u+p^{0}$
@@
Set the parameter values of Algorithm 2 as $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{19}{44\max\left\{1+\kappa_{f},6\right\}}}\right\}$, $\eta=\frac{1}{4\tau\max\left\{L_{f}+\mu_{f},6\mu_{f}\right\}}$, $\theta=\frac{15}{19\eta}$ and $\alpha=\frac{\mu_{f}}{4}$. Denote by $x^{*}$ the solution of problem (1). Then, for every $\varepsilon>0$, Algorithm 2 finds $x^{k}$ for which $\|x^{k}-x^{*}\|^{2}\leq\varepsilon$ using $O(\sqrt{\kappa_{f}}\log(1/\varepsilon))$ objective’s gradient computations, $O(\sqrt{\kappa_{f}}\sqrt{\kappa_{\mathbf{A}}}\log(1/\varepsilon))$ multiplications by $\mathbf{A}$ and $\mathbf{A}^{\top}$, and $O(\sqrt{\kappa_{f}}\sqrt{\kappa_{\mathbf{A}}}\sqrt{\kappa_{\mathbf{W}}}\log(1/\varepsilon))$ communication rounds (multiplications by $\mathbf{W}$).
@@ 5.2 Lower bounds
Below we describe how the sets $\mathcal{M}_{i}(k),\mathcal{H}_{i}(k)$ are updated.
@@
A proof of Theorem 2 is provided in Appendix B.
@@ 6 Experiments
where we randomly generate matrices $C_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, $\mathbf{A}_{i}\in\mathbb{R}^{m\times d_{i}}$ and vectors $d_{i}\in\mathbb{R}^{d_{i}}$, $b_{i}\in\mathbb{R}^{m}$ from the standard normal distribution. Local variables $x_{i}\in\mathbb{R}^{d_{i}}$ have the same dimension $d_{i}$, equal for all devices. Regularization parameter $\theta$ is $10^{-3}$. In the Fig. 1 we demonstrate the performance of the our method on the problem, that has the following parameters: $\kappa_{f}=3140$, $\kappa_{\mathbf{A}}=27$, $\kappa_{\mathbf{W}}=89$. There we use Erdős–Rényi graph topology with $n=20$ nodes. Local variables dimension is $d_{i}=3$ and number of linear constraints is $m=10$. We compare performance of Algorithm 2 with Tracking-ADMM algorithm Falsone et al. (2020) and DPMM algorithm Gong and Zhang (2023). Note that Tracking-ADMM and DPMM are proximal algorithms that solve a subproblem at each iteration. The choice of objective function in our simulations (linear regression) makes the corresponding proximal operator effectively computable via Conjugate Gradient algorithm Nesterov (2004) that uses gradient computations. Therefore, we measure the computational complexity of these methods in the number of gradient computations, not the number of proximal operator computations.
@@
Our algorithm exhibits the best convergence rates, as evidenced by the steepest slopes. The slopes vary for gradient calls, matrix multiplications, and communications. This is due to the fact that Algorithm 2 involves many communications per iteration, in contrast to DPMM and Tracking-ADMM, which make numerous gradient calls per iteration.
@@
Appendix A Missing proofs from Section 4
A.1 Proof of Lemma 1
@@
=$\overset{(d)}{=}\frac{\mu_{f}}{4}\|x^{\prime}-x\|^{2}+\frac{\mu_{f}\gamma^{2}\mu_{\mathbf{W^{\prime}}}}{8L_{\mathbf{A}}}\|y^{\prime}-y\|^{2},$
@@
where (a) is due to Assumption 1; (b) is due to Young’s inequality; (c) is due to Assumption 2, $y^{\prime}-y\in\mathcal{L}_{m}^{\perp}$, eq. 9 and eq. 6; (d) and (e) is due to eq. 13.
@@
=$\overset{(d)}{=}\frac{L_{f}+\mu_{f}}{2}\|x^{\prime}-x\|^{2}+\frac{\mu_{f}\gamma^{2}L_{\mathbf{W^{\prime}}}}{2L_{\mathbf{A}}}\|y^{\prime}-y\|^{2},$
@@
where (a) is due to Assumption 1; (b) is due to Young’s inequality; (c) is due to Assumption 2 and eq. 6; (d) and (e) is due to eq. 13. ∎
A.2 Proof of Lemma 2
@@
=$\overset{(b)}{=}\left(L_{\mathbf{A}}+(L_{\mathbf{A}}+\mu_{\mathbf{A}})\frac{L_{\mathbf{W^{\prime}}}}{\mu_{\mathbf{W^{\prime}}}}\right)\|z\|^{2},$
where (a) is due to Assumption 2 and eq. 6; (b) is due to eq. 13.
To derive the formula for $\mu_{\mathbf{B}}$, first of all, note that by eq. 9
@@
$\|\mathbf{B}^{\top}z\|^{2}$ =$\overset{(a)}{=}\|\mathbf{A}^{\top}(u+v)\|^{2}+\|\gamma\mathbf{W^{\prime}}(u+v)\|^{2}$
=$\overset{(b)}{=}\|\mathbf{A}^{\top}(u+v)\|^{2}+\|\gamma\mathbf{W^{\prime}}u\|^{2}$
@@
=$\overset{(e)}{=}-\|\mathbf{A}^{\top}u\|^{2}+\frac{1}{2}\langle v_{0},n\mathbf{S}v_{0}\rangle+\gamma^{2}\mu_{\mathbf{W^{\prime}}}\|u\|^{2}$
@@
=$\overset{(g)}{=}\frac{\mu_{\mathbf{A}}}{2}\|v\|^{2}+\mu_{\mathbf{A}}\|u\|^{2}$
@@
where (a) and (h) is due to the definitions of $u$ and $v$; (b) is due to the fact that $v\in\mathcal{L}_{m}$; (c) is due to eq. 6 and eq. 9; (d) uses Young’s inequality; (e) is due to the definitions of $v$ and $\mathbf{S}$, and $\|\mathbf{A}^{\top}v\|^{2}=\left\|\begin{pmatrix}\mathbf{A}_{1}^{\top}v_{0}\\ \vdots\\ \mathbf{A}_{n}^{\top}v_{0}\end{pmatrix}\right\|^{2}=\sum_{i=1}^{n}\|\mathbf{A}_{i}^{\top}v_{0}\|^{2}=\langle v_{0},\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{A}_{i}^{\top}v_{0}\rangle=\langle v_{0},n\mathbf{S}v_{0}\rangle$; (f) is due to Assumption 2 and the definition of $v$; (g) is due to eq. 13.
A.3 Proof of Theorem 1
@@
Let $\mathbf{M}$ be a matrix with $\mu_{\mathbf{M}}>0$, $\mathbf{r}\in\operatorname{\mathrm{range}}\mathbf{M}$ and $\mathbf{M}v_{0}=\mathbf{r}$. Then $\mathcal{P}_{\mathbf{M}}(\mathbf{M}^{\top}\mathbf{M})(v-v_{0})=v-\mathrm{\mathbf{Chebyshev}}(v,\mathbf{M},\mathbf{r})$, where $\mathrm{\mathbf{Chebyshev}}$ is defined as Algorithm 6.
@@
2:$n\coloneqq\left\lceil\sqrt{\frac{L_{\mathbf{M}}}{\mu_{\mathbf{M}}}}\right\rceil$
3:$\rho\coloneqq\big{(}L_{\mathbf{M}}-\mu_{\mathbf{M}}\big{)}^{2}/16$, $\nu\coloneqq(L_{\mathbf{M}}+\mu_{\mathbf{M}})/2$
4:$\delta^{0}\coloneqq-\nu/2$
5:$p^{0}\coloneqq-{\mathbf{M}}^{\top}({\mathbf{M}}v-\mathbf{r})/\nu$
6:$v^{1}\coloneqq v+p^{0}$
@@
Applying Lemma 3 to $\mathbf{W}$ and $\mathbf{B}^{\top}\mathbf{B}$, we derive that, due to eq. 19, it holds
@@
and by eq. 10 the polynomial $\mathcal{P}_{\mathbf{W}}$ has a degree of $\lceil\sqrt{\kappa}_{\mathbf{W}}\rceil$. Similarly, due to eq. 20, it holds
@@
We implement computation of the term $\mathbf{K}^{\top}(\mathbf{K}u-\mathbf{b^{\prime}})$ in line 6 of Algorithm 1 via Algorithm 5 by Lemma 5:
@@
Similarly, utilizing Lemma 5, we get
@@
where $\mathrm{\mathbf{mul}\mathbf{W^{\prime}}}$ is defined as Algorithm 3.
Therefore, Algorithm 2 is equivalent to Algorithm 1.
From eqs. 14 and 28, $\frac{\mu_{\mathbf{A}}+L_{\mathbf{A}}}{L_{\mathbf{A}}}\leq 2$ and $(19/11)^{2}\leq 3$, we get
@@
and substituting eqs. 34 and 33 to Proposition 1, we obtain as its direct corollary that $k=O(\sqrt{\kappa_{f}}\log(1/\varepsilon))$. Each iteration of Algorithm 2 require $O(1)$ computations of $\nabla F$, $O(\sqrt{\kappa_{\mathbf{B}}})=O(\sqrt{\kappa_{\mathbf{A}}})$ multiplications by $\mathbf{A},\mathbf{A}^{\top}$ and $O(\sqrt{\kappa_{\mathbf{A}}}\sqrt{\kappa_{\mathbf{W}}})$ multiplications by $\mathbf{W}$, which gives us the statement of Theorem 1. The values of the parameters $\tau,\eta,\theta,\alpha$ in Theorem 1 are derived from Proposition 1 as follows. We have $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{\kappa_{\mathbf{K}}}{\kappa_{G}}}\right\}=\min\left\{1,\frac{1}{2}\sqrt{\frac{19}{44\max\left\{1+\kappa_{f},6\right\}}}\right\}$ due to eqs. 35 and 33; $\eta=\frac{1}{4\tau L_{G}}=\frac{1}{4\tau\max\left\{L_{f}+\mu_{f},6\mu_{f}\right\}}$ due to eq. 31; $\theta=\frac{15}{19\eta}$ due to eq. 29 and $\alpha=\mu_{G}=\frac{\mu_{f}}{4}$ due to eq. 32. ∎
Appendix B Proof of Theorem 2
B.1 Dual problem
Let us construct the lower bound for the problem dual to the initial one. Consider primal problem with zero r.h.s. in constraints.
$\min_{x_{1},\ldots,x_{n}\in\mathbb{R}^{d}}~{}$ $\sum_{i=1}^{n}f_{i}(x_{i})$
s.t. $\sum_{i=1}^{n}\mathbf{A}_{i}x_{i}=0$

### v4

@@ 5.1 Algorithm
As stated in Lemma˜4, problem (20) is equivalent to problem (1). Due to Lemma˜1, its objective is strongly convex, allowing us to apply Algorithm˜1 to it. Using Lemma˜3, we obtain that the condition numbers of $\mathbf{W^{\prime}}$ and $\mathbf{K}$ are bounded as $O(1)$, but a single multiplication by $\mathbf{W^{\prime}}$ and $\mathbf{K},\mathbf{K}^{\top}$ translates to $O(\sqrt{\kappa_{\mathbf{W}}})$ multiplications by $\mathbf{W}$ and $O(\sqrt{\kappa_{\mathbf{B}}})$ multiplications by $\mathbf{B},\mathbf{B}^{\top}$ respectively.
We implement multiplications by $\mathbf{W^{\prime}}$ and $\mathbf{K},\mathbf{K}^{\top}$ through numerically stable Chebyshev iteration procedures given in Algorithms˜3 and 5, which only use decentralized communications and multiplications by $\mathbf{A},\mathbf{A}^{\top}$. Lemmas˜1, 2 and 3 allow us to express the complexity of Algorithm˜1 in terms of the parameters of the initial problem given in Assumptions˜1, 2 and 3. All this leads us to the following Theorem˜1, a detailed proof of which is provided in Appendix B.3, as well as the derivation of Algorithms˜3 and 5 and values of the parameters of Algorithm˜2.
@@
[…]
@@
[…]
@@
[…]
@@
Set the parameter values of Algorithm˜2 as $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{19}{44\max\left\{1+\kappa_{f},6\right\}}}\right\}$, $\eta=\frac{1}{4\tau\max\left\{L_{f}+\mu_{f},6\mu_{f}\right\}}$, $\theta=\frac{15}{19\eta}$ and $\alpha=\frac{\mu_{f}}{4}$. Denote by $x^{*}$ the solution of problem (1). Then, for every $\varepsilon>0$, Algorithm˜2 finds $x^{k}$ for which $\|x^{k}-x^{*}\|^{2}\leq\varepsilon$ using $O(\sqrt{\kappa_{f}}\log(1/\varepsilon))$ objective’s gradient computations, $O(\sqrt{\kappa_{f}}\sqrt{\kappa_{\mathbf{A}}}\log(1/\varepsilon))$ multiplications by $\mathbf{A}$ and $\mathbf{A}^{\top}$, and $O(\sqrt{\kappa_{f}}\sqrt{\kappa_{\mathbf{A}}}\sqrt{\kappa_{\mathbf{W}}}\log(1/\varepsilon))$ communication rounds (multiplications by $\mathbf{W}$).
@@ 5.2 Lower bounds
Memory is initialized as $\mathcal{M}_{i}(0)=\left\{0\right\},~\mathcal{H}_{i}(0)=\left\{0\right\}$. Below we describe how the sets $\mathcal{M}_{i}(k),\mathcal{H}_{i}(k)$ are updated.
@@
A proof of Theorem 2 is provided in Appendix C.
@@ 6 Experiments
where we randomly generate matrices $C_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, $\mathbf{A}_{i}\in\mathbb{R}^{m\times d_{i}}$ and vectors $d_{i}\in\mathbb{R}^{d_{i}}$, $b_{i}\in\mathbb{R}^{m}$ from the standard normal distribution. Local variables $x_{i}\in\mathbb{R}^{d_{i}}$ have the same dimension $d_{i}$, equal for all devices. Regularization parameter $\theta$ is $10^{-3}$. In the Fig. 1 we demonstrate the performance of the our method on the problem, that has the following parameters: $\kappa_{f}=3140$, $\kappa_{\mathbf{A}}=27$, $\kappa_{\mathbf{W}}=89$. There we use Erdős–Rényi graph topology with $n=20$ nodes. Local variables dimension is $d_{i}=3$ and number of linear constraints is $m=10$. We compare performance of Algorithm˜2 with Tracking-ADMM algorithm Falsone et al. (2020) and DPMM algorithm Gong and Zhang (2023). Note that Tracking-ADMM and DPMM are proximal algorithms that solve a subproblem at each iteration. The choice of objective function in our simulations (linear regression) makes the corresponding proximal operator effectively computable via Conjugate Gradient algorithm Nesterov (2004) that uses gradient computations. Therefore, we measure the computational complexity of these methods in the number of gradient computations, not the number of proximal operator computations.
@@
Our algorithm exhibits the best convergence rates, as evidenced by the steepest slopes. The slopes vary for gradient calls, matrix multiplications, and communications. This is due to the fact that Algorithm˜2 involves many communications per iteration, in contrast to DPMM and Tracking-ADMM, which make numerous gradient calls per iteration.
7 Acknowledgements
The work was supported by MIPT based Center of National Technology Initiatives in the field of Artificial Intelligence for the purposes of the "road map" of Artificial Intelligence development up to 2030 and supported by NTI Foundation (agreement No.70-2021-00207 dated 22.11.2021, identifier  [^00] S507521QYL0002).
@@ 7 Acknowledgements
Appendix A Reduction of consensus optimization to coupled constraints
As mentioned in Section˜1, the consensus optimization problem
$\min_{x_{1},\ldots,x_{n}\in\mathbb{R}^{d}}\sum_{i=1}^{n}f_{i}(x_{i})\quad\text{s.t.}\quad x_{1}=x_{2}=\ldots=x_{n}$
can be reduced to the problem with coupled constraints (1). During the review process, we were asked whether the complexity of Algorithm˜2 could be reduced to match that of optimal algorithms for consensus optimization Scaman et al. [2017].
It turns out that, for a general-purpose first-order decentralized algorithm for problems with coupled constraints (as defined in Section˜5.2), the communication complexity is worse by at least a factor of $\sqrt{n}$. This conclusion is based on our complexity lower bounds in Theorem˜2 and the following lower bound on $\kappa_{\mathbf{A}}$ for any suitable matrix $\mathbf{A}$.
Let $d=1$. Consider an arbitrary horizontal-block matrix $\mathbf{A}^{\prime}=(\mathbf{A}_{1}\ldots\mathbf{A}_{n})\in\mathbb{R}^{m\times n}$ such that $\ker\mathbf{A}^{\prime}=\mathcal{L}_{1}=\mathrm{Span}\{\vec{1}_{n}\}$, which corresponds to the consensus constraint. By definition (4), we have
$\mu_{\mathbf{A}}=\frac{1}{n}\lambda_{\min^{+}}\left(\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\right)=\frac{1}{n}\lambda_{\min^{+}}\left(\mathbf{A}^{\prime}\mathbf{A}^{\prime\top}\right)=\frac{1}{n}\sigma_{\min^{+}}^{2}(\mathbf{A}^{\prime}),$
and
$L_{\mathbf{A}}=\max_{i=1\ldots n}\sigma^{2}_{\max}(\mathbf{A}_{i})=\|\mathbf{A}_{i^{*}}\|^{2},$
where $i^{*}$ denotes the index of the column achieving the maximum. We can upper bound the minimal positive singular value using a vector $v$, whose components are all zero except for one $1$ and one 1-1, implying $\|v\|^{2}=2$. Since the sum of its components is zero, $v$ is orthogonal to $\ker\mathbf{A}^{\prime}$, thus
$\sigma_{\min^{+}}^{2}(\mathbf{A}^{\prime})=\min_{\begin{subarray}{c}v:~\|v\|>0,\\ v\bot\vec{1}\end{subarray}}\|\mathbf{A}^{\prime}v\|^{2}/\|v\|^{2}\leq\|\mathbf{A}_{i}-\mathbf{A}_{j}\|^{2}/2\leq 2\|\mathbf{A}_{i^{*}}\|^{2}.$
It follows that
$\kappa_{\mathbf{A}}=\frac{L_{\mathbf{A}}}{\mu_{\mathbf{A}}}\geq\frac{\|\mathbf{A}_{i^{*}}\|^{2}}{\frac{2}{n}\|\mathbf{A}_{i^{*}}\|^{2}}=\frac{n}{2}.$
This bound is nearly tight, because we can achieve $\kappa_{\mathbf{A}}=n-1$ by setting $\mathbf{A}^{\prime}$ to be the Laplacian matrix or the incidence matrix of a complete graph on $n$ nodes.
Substituting this into the complexity lower bounds in Theorem˜2, we find that the number of communication rounds is lower bounded by $\Omega\left(\sqrt{\kappa_{f}}\sqrt{n}\sqrt{\kappa_{\mathbf{W}}}\log\left(\frac{1}{\varepsilon}\right)\right)$ for any choice of $\mathbf{A}$, while the optimal complexity for first order decentralized consensus optimization is $O\left(\sqrt{\kappa_{f}}\sqrt{\kappa_{\mathbf{W}}}\log\left(\frac{1}{\varepsilon}\right)\right)$.
Appendix B Missing proofs from Section 4
B.1 Proof of Lemma 1
@@
$\overset{(d)}{=}\frac{\mu_{f}}{4}\|x^{\prime}-x\|^{2}+\frac{\mu_{f}\gamma^{2}\mu_{\mathbf{W^{\prime}}}}{8L_{\mathbf{A}}}\|y^{\prime}-y\|^{2},$
@@
where (a) is due to Assumption˜1; (b) is due to Young’s inequality; (c) is due to Assumption˜2, $y^{\prime}-y\in\mathcal{L}_{m}^{\perp}$, eq.˜8 and eq.˜5; (d) and (e) is due to eq.˜12.
@@
$\overset{(d)}{=}\frac{L_{f}+\mu_{f}}{2}\|x^{\prime}-x\|^{2}+\frac{\mu_{f}\gamma^{2}L_{\mathbf{W^{\prime}}}}{2L_{\mathbf{A}}}\|y^{\prime}-y\|^{2},$
@@
where (a) is due to Assumption˜1; (b) is due to Young’s inequality; (c) is due to Assumption˜2 and eq.˜5; (d) and (e) is due to eq.˜12. ∎
B.2 Proof of Lemma 2
@@
$\overset{(b)}{=}\left(L_{\mathbf{A}}+(L_{\mathbf{A}}+\mu_{\mathbf{A}})\frac{L_{\mathbf{W^{\prime}}}}{\mu_{\mathbf{W^{\prime}}}}\right)\|z\|^{2},$
where (a) is due to Assumption˜2 and eq.˜5; (b) is due to eq.˜12.
To derive the formula for $\mu_{\mathbf{B}}$, first of all, note that by eq.˜8
@@
$\|\mathbf{B}^{\top}z\|^{2}$ $\overset{(a)}{=}\|\mathbf{A}^{\top}(u+v)\|^{2}+\|\gamma\mathbf{W^{\prime}}(u+v)\|^{2}$
$\overset{(b)}{=}\|\mathbf{A}^{\top}(u+v)\|^{2}+\|\gamma\mathbf{W^{\prime}}u\|^{2}$
@@
$\overset{(e)}{=}-\|\mathbf{A}^{\top}u\|^{2}+\frac{1}{2}\langle v_{0},n\mathbf{S}v_{0}\rangle+\gamma^{2}\mu_{\mathbf{W^{\prime}}}\|u\|^{2}$
@@
$\overset{(g)}{=}\frac{\mu_{\mathbf{A}}}{2}\|v\|^{2}+\mu_{\mathbf{A}}\|u\|^{2}$
@@
where (a) and (h) is due to the definitions of $u$ and $v$; (b) is due to the fact that $v\in\mathcal{L}_{m}$; (c) is due to eq.˜5 and eq.˜8; (d) uses Young’s inequality; (e) is due to the definitions of $v$ and $\mathbf{S}$, and $\|\mathbf{A}^{\top}v\|^{2}=\left\|\begin{pmatrix}\mathbf{A}_{1}^{\top}v_{0}\\ \vdots\\ \mathbf{A}_{n}^{\top}v_{0}\end{pmatrix}\right\|^{2}=\sum_{i=1}^{n}\|\mathbf{A}_{i}^{\top}v_{0}\|^{2}=\langle v_{0},\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{A}_{i}^{\top}v_{0}\rangle=\langle v_{0},n\mathbf{S}v_{0}\rangle$; (f) is due to Assumption˜2 and the definition of $v$; (g) is due to eq.˜12.
B.3 Proof of Theorem 1
@@
Let $\mathbf{M}$ be a matrix with $\mu_{\mathbf{M}}>0$, $\mathbf{r}\in\operatorname{\mathrm{range}}\mathbf{M}$ and $\mathbf{M}v_{0}=\mathbf{r}$. Then $\mathcal{P}_{\mathbf{M}}(\mathbf{M}^{\top}\mathbf{M})(v-v_{0})=v-\mathrm{\mathbf{Chebyshev}}(v,\mathbf{M},\mathbf{r})$, where $\mathrm{\mathbf{Chebyshev}}$ is defined as Algorithm˜6.
@@
[…]
@@
Applying Lemma˜3 to $\mathbf{W}$ and $\mathbf{B}^{\top}\mathbf{B}$, we derive that, due to eq.˜18, it holds
@@
and by eq.˜6 the polynomial $\mathcal{P}_{\mathbf{W}}$ has a degree of $\lceil\sqrt{\kappa}_{\mathbf{W}}\rceil$. Similarly, due to eq.˜19, it holds
@@
We implement computation of the term $\mathbf{K}^{\top}(\mathbf{K}u-\mathbf{b^{\prime}})$ in line 6 of Algorithm˜1 via Algorithm˜5 by Lemma˜5:
@@
Similarly, utilizing Lemma˜5, we get
@@
where $\mathrm{\mathbf{mul}\mathbf{W^{\prime}}}$ is defined as Algorithm˜3.
Therefore, Algorithm˜2 is equivalent to Algorithm˜1.
From eqs.˜13 and 27, $\frac{\mu_{\mathbf{A}}+L_{\mathbf{A}}}{L_{\mathbf{A}}}\leq 2$ and $(19/11)^{2}\leq 3$, we get
@@
and substituting eqs.˜33 and 32 to Proposition˜1, we obtain as its direct corollary that $k=O(\sqrt{\kappa_{f}}\log(1/\varepsilon))$. Each iteration of Algorithm˜2 require $O(1)$ computations of $\nabla F$, $O(\sqrt{\kappa_{\mathbf{B}}})=O(\sqrt{\kappa_{\mathbf{A}}})$ multiplications by $\mathbf{A},\mathbf{A}^{\top}$ and $O(\sqrt{\kappa_{\mathbf{A}}}\sqrt{\kappa_{\mathbf{W}}})$ multiplications by $\mathbf{W}$, which gives us the statement of Theorem˜1. The values of the parameters $\tau,\eta,\theta,\alpha$ in Theorem˜1 are derived from Proposition˜1 as follows. We have $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{\kappa_{\mathbf{K}}}{\kappa_{G}}}\right\}=\min\left\{1,\frac{1}{2}\sqrt{\frac{19}{44\max\left\{1+\kappa_{f},6\right\}}}\right\}$ due to eqs.˜34 and 32; $\eta=\frac{1}{4\tau L_{G}}=\frac{1}{4\tau\max\left\{L_{f}+\mu_{f},6\mu_{f}\right\}}$ due to eq.˜30; $\theta=\frac{15}{19\eta}$ due to eq.˜28 and $\alpha=\mu_{G}=\frac{\mu_{f}}{4}$ due to eq.˜31. ∎
Appendix C Proof of Theorem 2
C.1 Dual problem
Let us construct the lower bound for the problem dual to the initial one. Consider primal problem with zero r.h.s. in constraints
$\min_{x_{1},\ldots,x_{n}\in\ell_{2}}$ $\sum_{i=1}^{n}f_{i}(x_{i})$ (35)
s.t. $\sum_{i=1}^{n}\mathbf{A}_{i}x_{i}=0.$